# Tyche: Stochastic In-Context Learning for Medical Image Segmentation

### v1

@@ B.3 Network for Tyche-IS
For the Tyche-IS network, we use the baseline UniverSeg [19], trained with the same data as Tyche-TS, same batch size, and same context size. At test time, we use the augmentations in Table 7. Figure 12 shows example augmentation on different samples.
@@
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.25 contrast$\in[0.5,1.5]$
Table 7: Augmentations used for Tyche-IS. We focus on intensity transforms, to avoid inverting the prediction. For each image, an augmentation is sampled with probability $p$.
@@ B.4 Benchmarks
(b) Light Augmentation. Gaussian bur, Gaussian noise, elastic transforms are each applied to the target with probability $p$.
@@
translate$\in[0,0.2]$
Random Affine 0.5 scale$\in[0.8,1.1]$
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast$\in[0.5,1.5]$
@@ D.6 Augmentations
brightness$\in[-0.1,0.1]$,
Brightness Contrast 0.25 contrast$\in[0.5,1.5]$
Table 9: High augmentations used to validate Tyche-IS. We focus on intensity transforms, to avoid inverting the prediction. For each image, an augmentation is sampled with probability $p$.
@@ E.1 Performances on other Metrics
Evaluating the quality of different predictions can be challenging especially when different annotators are available. We proposed best candidate Dice score and Generalized Energy Distance. Some also analysed sample diversity [95], and Hungarian Matching [68, 133]. Sample diversity consists in measuring the agreement between candidate predictions $\hat{\mathcal{Y}}$, rewarding most diverse sets of candidates:
@@
In the context of stochastic predictions, Hungarian Matching [70] consists in matching the set of predictions with the set of annotations, so that an overall metric is minimized. We use negative Dice score. One limitation of this method is that it has to be adapted when the number of annotators does not match the number of prediction. The most widely used fix is to artificially inflate the number of predictions and the number of annotations to reach the least common multiple [68]. We use this strategy here as well.
@@ E.3 Additional Visualizations
ACDC [14]
Left and right ventricular endocardium
99
cine-MRI
AMOS [54]
Abdominal organ segmentation
240
CT, MRI
BBBC003 [83]
Mouse embryos
15
Microscopy
BBBC038 [20]
Nuclei images
670
Microscopy
BUID [2]
Breast tumors
647
Ultrasound
BrainDev. [38, 37, 72, 114]
Adult and Neonatal Brain Atlases
53
multi-modal MRI
BRATS [7, 8, 93]
Brain tumors
6,096
multi-modal MRI
BTCV [75]
Abdominal Organs
30
CT
BUS [136]
Breast tumor
163
Ultrasound
CAMUS [77]
Four-chamber and Apical two-chamber heart
500
Ultrasound
CDemris [56]
Human Left Atrial Wall
60
CMR
CHAOS [60, 58]
Abdominal organs (liver, kidneys, spleen)
40
CT, T2-weighted MRI
CheXplanation [110]
Chest X-Ray observations
170
X-Ray
CT-ORG[106]
Abdominal organ segmentation (overlap with LiTS)
140
CT
DRIVE [122]
Blood vessels in retinal images
20
Optical camera
EOphtha [27]
Eye Microaneurysms and Diabetic Retinopathy
102
Optical camera
FeTA [102]
Fetal brain structures
80
Fetal MRI
FetoPlac [10]
Placenta vessel
6
Fetoscopic optical camera
HMC-QU [28, 65]
4-chamber (A4C) and apical 2-chamber (A2C) left wall
292
Ultrasound
HipXRay [40]
Ilium and femur
140
X-Ray
I2CVB [78]
Prostate (peripheral zone, central gland)
19
T2-weighted MRI
IDRID [103]
Diabetic Retinopathy
54
Optical camera
ISLES [43]
Ischemic stroke lesion
180
multi-modal MRI
KiTS [42]
Kidney and kidney tumor
210
CT
LGGFlair [18, 91]
TCIA lower-grade glioma brain tumor
110
MRI
LiTS [16]
Liver Tumor
131
CT
LUNA [115]
Lungs
888
CT
MCIC [36]
Multi-site Brain regions of Schizophrenic patients
390
T1-weighted MRI
MSD [118]
Collection of 10 Medical Segmentation Datasets
3,225
CT, multi-modal MRI
NCI-ISBI [17]
Prostate
30
T2-weighted MRI
OASIS [46, 88]
Brain anatomy
414
T1-weighted MRI
OCTA500 [79]
Retinal vascular
500
OCT/OCTA
PanDental [1]
Mandible and Teeth
215
X-Ray
PAXRay [112]
Thoracic organs
880
X-Ray
PROMISE12 [82]
Prostate
37
T2-weighted MRI
PPMI [89]
Brain regions of Parkinson patients
1,130
T1-weighted MRI
ROSE [86]
Retinal vessel
117
OCT/OCTA
SCD [104]
Sunnybrook Cardiac Multi-Dataset Collection
100
cine-MRI
SegTHOR [74]
Thoracic organs (heart, trachea, esophagus)
40
CT
SpineWeb [138]
Vertebrae
15
T2-weighted MRI
ToothSeg [50]
Individual teeth
598
X-Ray
WBC [139]
White blood cell and nucleus
400
Microscopy
WMH [71]
White matter hyper-intensities
60
multi-modal MRI
WORD [85]
Organ segmentation
120
CT
@@
LIDC-IDRI [4]
Lung Nodules
1018
CT
QUBIQ [92]
Brain, kidney, pancreas and prostate
209
MRI T1, Multimodal MRI, CT
STARE [47]
Blood vessels in retinal images
20
Optical camera

### v2

@@ B.3 Network for Tyche-IS
For the Tyche-IS network, we use the baseline UniverSeg [butoi2023universeg], trained with the same data as Tyche-TS, same batch size, and same context size. At test time, we use the augmentations in Table 7. Figure 12 shows example augmentation on different samples.
@@
brightness[$\in[-0.1,0.1]$,
Brightness Contrast 0.25 contrast[$\in[0.5,1.5]$
Table 7: Augmentations used for Tyche-IS. We focus on intensity transforms, to avoid inverting the prediction. For each image, an augmentation is sampled with probability pp.
@@ B.4 Benchmarks
(b) Light Augmentation. Gaussian bur, Gaussian noise, elastic transforms are each applied to the target with probability pp.
@@
translate[$\in[0,0.2]$
Random Affine 0.5 scale[$\in[0.8,1.1]$
brightness[$\in[-0.1,0.1]$,
Brightness Contrast 0.5 contrast[$\in[0.5,1.5]$
@@ D.6 Augmentations
brightness[$\in[-0.1,0.1]$,
Brightness Contrast 0.25 contrast[$\in[0.5,1.5]$
Table 9: High augmentations used to validate Tyche-IS. We focus on intensity transforms, to avoid inverting the prediction. For each image, an augmentation is sampled with probability pp.
@@ E.1 Performances on other Metrics
Evaluating the quality of different predictions can be challenging especially when different annotators are available. We proposed best candidate Dice score and Generalized Energy Distance. Some also analysed sample diversity [monteiro2020stochastic], and Hungarian Matching [kohl2019hierarchical, zbinden2023stochastic]. Sample diversity consists in measuring the agreement between candidate predictions $\hat{\mathcal{Y}}$, rewarding most diverse sets of candidates:
@@
In the context of stochastic predictions, Hungarian Matching [kuhn1955hungarian] consists in matching the set of predictions with the set of annotations, so that an overall metric is minimized. We use negative Dice score. One limitation of this method is that it has to be adapted when the number of annotators does not match the number of prediction. The most widely used fix is to artificially inflate the number of predictions and the number of annotations to reach the least common multiple [kohl2019hierarchical]. We use this strategy here as well.
@@ E.3 Additional Visualizations
ACDC [ACDC] Left and right ventricular endocardium 99 cine-MRI
AMOS [AMOS] Abdominal organ segmentation 240 CT, MRI
BBBC003 [BBBC003] Mouse embryos 15 Microscopy
BBBC038 [BBBC038] Nuclei images 670 Microscopy
BUID [BUID] Breast tumors 647 Ultrasound
BrainDev. [gousias2012magnetic, BrainDevFetal, BrainDevelopment, serag2012construction] Adult and Neonatal Brain Atlases 53 multi-modal MRI
BRATS [BRATS, bakas2017advancing, menze2014multimodal] Brain tumors 6,096 multi-modal MRI
BTCV [BTCV] Abdominal Organs 30 CT
BUS [Bus] Breast tumor 163 Ultrasound
CAMUS [CAMUS] Four-chamber and Apical two-chamber heart 500 Ultrasound
CDemris [cDemris] Human Left Atrial Wall 60 CMR
CHAOS [Chaos_1, Chaos_2] Abdominal organs (liver, kidneys, spleen) 40 CT, T2-weighted MRI
CheXplanation [CheXplanation] Chest X-Ray observations 170 X-Ray
CT-ORG[CT_ORG] Abdominal organ segmentation (overlap with LiTS) 140 CT
DRIVE [DRIVE] Blood vessels in retinal images 20 Optical camera
EOphtha [EOphtha] Eye Microaneurysms and Diabetic Retinopathy 102 Optical camera
FeTA [FeTA] Fetal brain structures 80 Fetal MRI
FetoPlac [FetoPlac] Placenta vessel 6 Fetoscopic optical camera
HMC-QU [HMC-QU, kiranyaz2020left] 4-chamber (A4C) and apical 2-chamber (A2C) left wall 292 Ultrasound
HipXRay [HipXRay] Ilium and femur 140 X-Ray
I2CVB [I2CVB] Prostate (peripheral zone, central gland) 19 T2-weighted MRI
IDRID [IDRID] Diabetic Retinopathy 54 Optical camera
ISLES [ISLES] Ischemic stroke lesion 180 multi-modal MRI
KiTS [KiTS] Kidney and kidney tumor 210 CT
LGGFlair [buda2019association, LGGFlair] TCIA lower-grade glioma brain tumor 110 MRI
LiTS [LiTS] Liver Tumor 131 CT
LUNA [LUNA] Lungs 888 CT
MCIC [MCIC] Multi-site Brain regions of Schizophrenic patients 390 T1-weighted MRI
MSD [MSD] Collection of 10 Medical Segmentation Datasets 3,225 CT, multi-modal MRI
NCI-ISBI [NCI-ISBI] Prostate 30 T2-weighted MRI
OASIS [OASIS-proc, OASIS-data] Brain anatomy 414 T1-weighted MRI
OCTA500 [OCTA500] Retinal vascular 500 OCT/OCTA
PanDental [PanDental] Mandible and Teeth 215 X-Ray
PAXRay [PAXRay] Thoracic organs 880 X-Ray
PROMISE12 [Promise12] Prostate 37 T2-weighted MRI
PPMI [PPMI] Brain regions of Parkinson patients 1,130 T1-weighted MRI
ROSE [Rose] Retinal vessel 117 OCT/OCTA
SCD [SCD] Sunnybrook Cardiac Multi-Dataset Collection 100 cine-MRI
SegTHOR [SegTHOR] Thoracic organs (heart, trachea, esophagus) 40 CT
SpineWeb [SpineWeb] Vertebrae 15 T2-weighted MRI
ToothSeg [ToothSeg] Individual teeth 598 X-Ray
WBC [WBC] White blood cell and nucleus 400 Microscopy
WMH [WMH] White matter hyper-intensities 60 multi-modal MRI
WORD [Word] Organ segmentation 120 CT
@@
LIDC-IDRI [armato2011lung] Lung Nodules 1018 CT
QUBIQ [qubiq] Brain, kidney, pancreas and prostate 209 MRI T1, Multimodal MRI, CT
STARE [STARE] Blood vessels in retinal images 20 Optical camera